### v1

@@ 3.1.2 Software
Software was also required. Server and client programs were written to send LiDAR, RGB, and thermal data, as well as control signals, between the gondola and an external laptop. TigerVNC was used to remotely develop the software on the Raspberry Pi.303030https://tigervnc.org/ As well, we built a user interface with color-coded functions in Python and OpenCV,313131https://opencv.org/ shown in Fig. 3. While we created Arduino code to control the blimp in various ways, to avoid potential problems if communication broke down, we ended up using commands that initiate short bursts rather than continuous motion.
One observation when developing software was how little information we could find on how to set up a reasonably realistic mock-up crime scene for testing; e.g., what kind of evidence is typical and where it is usually found. In addition to the challenge that crime classification schemes differ regionally, in line with the diversity of criminal law systems, Google search hits for random crime scene generation seemed to bring up text summaries of crimes intended to help stimulate writers (e.g., "Murder Scene Generator"323232https://perchance.org/znevive9bn). We could also not find any datasets online of crime scene sketches. One reason for the lack of crime data could be due to a view that such data are sensitive and private, for victims and their families, and could also spark copy-cat crimes. While one or two mock-up sketches were available (e.g., Galanakis et al. made available a scan of one mock-up crime scene [30], and our partners at the Swedish Police also shared a scan of a mock-up crime scene at NFC Villa), as the saying goes, no two crime scenes are the same. Thus, to get started, we built two simplified programs to simulate how evidence might be distributed at a crime scene, as shown in Fig. 4. The first general program allows a user to select a floor plan and a kind of crime (homicide, assault, burglary, or arson), then randomly computes which kinds of evidence are found and where. For initial floor plans, we used (1) a LiDAR scan of the National Forensics Center (NFC) villa, used by our partners at the Swedish Police to train, and (2) a virtual depiction of Halmstad Intelligent Home (HINT), a 50 m2 mock-up sensorized living environment at our university. All probability parameters can be assigned or tweaked by the user. The second program, intended for initial testing, simply creates a single "heap" of evidence, taking a list of evidence and using it to fill a grid. Large evidence comprising more than one cell is filled in first, then smaller evidence.
@@ 3.2 Exploring Possibilities for Indoor CSA Blimps
With the developed platform, based on the literature, we speculated that it could be interesting to check how much wind the blimp creates, how easy it is to control manually, how recordings of evidence can be further processed into 3D maps, and if thermal traces could be used. Furthermore, looking toward the future, we wished to check automatic detection, classification of bloodstains, and how a blimp could plan itself how to move around to acquire data. For evidence to use in testing, we gathered a photo of a firearm, two knives (one large and metallic, and one small and plastic), a pair of shoes, and three sheets with mock-up blood (red carmine dye–E120) to simulate passive, active, and transfer stains.
@@ 3.2.1 Disturbance
To test the degree to which sensitive evidence might be disturbed by a blimp, we placed a Mastech MS6252A digital anemometer on the ground beside the mock-up evidence. As a result, the wind sensor reading was 0.0 m/s even when passing the blimp even quite close to the evidence ( 20 cm). For comparison, passing a typical small drone, the DJI Ryze Tello, over the evidence at 1.2 m resulted in a wind speed of 0.6 to 0.8 m/s, which notably caused most of our evidence to scatter (viz., the sheets of paper containing mock-up blood and the photo of a firearm).
@@ 3.2.2 Sensing while drifting
Next, we checked on the blimp’s ability to move and sense evidence. The blimp was controlled manually to obtain 10 data samples. In each trial, a random ordering of the evidence was calculated using our Python program, then the physical evidence was manually placed by a human. As a result, on average, the blimp captured 5.2 of 7 objects (74.3%) in its view. The reason for missing some evidence sometimes was because the blimp was susceptible to drafts and small imbalances that sometimes caused it to veer.
@@
For easy viewing, we wished to explore converting video from the blimp to a 3D model. To test the waters, a video from the blimp was passed to the free version of 3DF Zephyr to create a 3D textured mesh. The result is shown in the Fig. 5(a). Some rough areas existed, e.g., on the mannequin "corpse" to the left-hand side of the image), since the free version of this software only uses the first 50 photos extracted from the video. (Due to this limitation, we also sped up the video to ensure the whole scene would be included.) However, we believe that the time required to generate a 3D map (approximately 3 minutes) was promising.
@@ 3.2.4 Thermal Sensing
In each trial controlling the blimp, we experimented with also touching a knife and the shoes directly beforehand. As a result, we noted that some warm areas could be seen in the thermal data, as in Fig. 5(b), but in general, the warmed objects (especially the smaller knife) seemed hard to distinguish due to our sensor’s low resolution. As well, a checkerboard pattern appears due to the blimp’s motion and the manner in which the MLX90640 sensor updates (in two passes, such that half of the pixels update slightly later). Thus, we believe that a higher resolution sensor with different update process would be required some day in the future for identifying recently-touched, non-metallic objects. (The MLX90640 could however still serve some other use, such as detecting investigators to follow, or hidden persons).
@@ 3.2.5 Object detection
It could save time if the blimp could be automated to perform important tasks like detection. This could then be highlighted in a 3D map in reports to investigators, or used to focus the blimp’s attention to more thoroughly investigate certain areas. As an initial check, we tried inputting an image taken by the blimp to a pretrained generic YOLO detection model (YOLOv8l)333333https://docs.ultralytics.com/ and ChatGPT.343434hhttps://chatgpt.com The first result in Fig. 5 (c) shows that while one knife is detected correctly, a bloody handprint is mistaken as a horse, and other objects are not detected; these errors could be due to low resolution or slightly dark lighting (approximately 300 lux in our lab when the image was taken). ChatGPT, however, correctly detected all objects. This seemed promising to us, and suggested using either an adapted version of YOLO with sufficient data for guns and blood, or a foundation model.
@@ 3.2.6 Bloodstain Classification
To explore classifying the three main kinds of bloodstains, we tried processing a photo of our evidence heap. First we color-picked red, then assessed the area and eccentricity of an ellipse fitted to detected contours of reasonable size. We estimated that transfer patterns from a hand might have high area, passive drops should have low eccentricity, and active drops high eccentricity, as in Fig. 5(d). This resulted in an classification accuracy of 74.1% (20 of 27 detected blood contours being correctly classified). Adding in false positives from color-picking (e.g., detected in the red tags on the shoes), the overall accuracy was 69.0% (20 of 29), which seemed promising for initial prototyping.
@@ 3.2.7 Coverage path planning for crime scenes
An automated blimp could move without piloting. Given that a full solution to this complex problem wouldn’t be possible within the scope of the current paper, we brainstormed to identify some variables of interest and formulate the general problem. We believe the blimp’s path planning strategy should depend on:
How precise blimp’s movements are–if imprecise, the blimp should stay farther away to avoid crashing, and more overlap might be required.
How accurate the blimp’s sensors are–if inaccurate, the blimp should stay closer to the evidence to achieve proper documentation.
How urgent the task is–if the blimp is needed elsewhere, a rapid path could be calculated; else in an important case, more time can be taken.
If the blimp should also gather evidence–then it should probably be close.
How much detail is desired–if the blimp should see under obstacles like tables and beds.
@@
In general, however, we believe that indoor blimp crime scene analysis is characterized by the following requirements:
Some speed is desired, given that evidence can become contaminated or degraded over time.
Some overlap is desired for 3D mapping; much (unneeded) overlap however would be undesirable since excess motion could increase the risk of disrupted evidence.
While human domestic environments can be highly complex (e.g., ceiling obstacles like lamps, fans, alarms, or chandeliers, and ground obstacles like chairs, clothes, etc. and L-shaped rooms or hallways and walk-in closets), rooms tend to often be rectangular.
The blimp should try to restrict turning and vertical changes to avoid using up its battery. For example, the blimp can fly over low objects (e.g., to see on tables, beds)–although, compared to an outdoors scene, height is more constrained.
People in the room might intentionally or unintentionally obstruct the robot.
The blimp could be acting as part of a swarm.
The environment might be known (e.g., from floor plans) or unknown.
Based on this, both simplified and more complex solutions seemed possible. For simplicity, we propose that a snaking algorithm combined with a random walk and/or wall-following could be a starting point for simple testing, that is intuitive and easy to implement and carry out. For maximum detail, the blimp could calculate a path where it flies just above surfaces; else for efficiency, the blimp could calculate a space below any ceiling obstacles and above any ground obstacles. As well, the blimp could scan from high-up first, then return to any areas that seem interesting (e.g., where evidence has been discovered). Alternatively, a deep RL algorithm could be used to document in a human-like way (e.g., based on data of human investigators): to achieve high efficiency (reduce flight time, avoiding redundant re-coverage), improve map quality (ensure good feature overlap and loop closure, while avoiding poor viewing angles, low lighting, occlusions, shadows or reflective surfaces), and be room-agnostic (function successfully in unknown or specially designed rooms or buildings, while avoiding hand-engineering edge cases), and robot-agnostic (different blimps could use the same process to learn how to move well). A cost might be more difficulty in the implementation (complex equations, much training data required, tuning), and loss of explainability.
@@ 4 Discussion
In summary, we explored a new design idea, of using a blimp to record sensitive evidence at an indoor crime scene. To our best knowledge, our new prototype is the first small indoor blimp to carry a thermal camera, or three LiDARs, or a minicomputer (rather than a microcontroller), etc., and various hardware challenges were discussed. Some software was also developed to simulate crime scenes. As a result of our exploration, we found some support that drones can disturb sensitive evidence and confirmed our premise that blimps can keep wind to a minimum. As well, manually piloting was possible but imperfect (sensing 74.3% of evidence). A 3D map was generated within a few minutes using free software. Some recommendations for thermal sensing were given. Toward automating crime scene analysis, all objects in the blimp’s camera view were detected correctly by ChatGPT. As well, the three main kinds of bloodstain were detected and classified with an accuracy of 74.1% via color-picking and consideration of contour area and eccentricity. Finally, we discussed path-planning considerations and strategies. As well, a video and code have been made available.353535https://youtu.be/5ep9UeChn68,363636https://github.com/martincooney/Crime_Blimp
@@ 4.1 Limitations and Future Work
The current study is limited by its exploratory design approach, working with mock-up evidence in a lab. The blimp was large, with component weights unoptimized; surface mount/low power components (like a Raspberry Pi Pico instead of Zero 2 W), or Dist-YOLO in place of LiDARs [31], to reduce weight and size. As well, we used a remote laptop, which, although common in the related studies noted, would not be secure at a real crime scene, given that criminals could find a way to hack (intercept or jam) communications. Our crime scene generating software is currently based on guesses and should be trained with real data to enable more realistic outputs. To help blimp operators, control algorithms should be developed. Other applications of heat sensing, and other miniature thermal cameras, should be tested. A more comprehensive study should be done on identifying and detecting the various kinds of evidence that could be important. For classifying bloodstains, only a single image was used. A reasonable number of (real) training and test data should be used in conjunction with a more advanced, machine-learning approach to achieve better results. As well, real composites of various kinds of bloodstains should be analyzed to infer the story of what happened, and further analysis can reveal, e.g., if a criminal is right or left-handed, the type of weapon used, what made a transfer stain (e.g., a hand or a shoe), which wounds were first, if death was immediate or not, and the number of blows, etc. Furthermore, path planning should be implemented and tested on our prototype.
As well, our focus was limited to sensitive crime evidence, whereas remote-monitoring blimps could be useful for other inspections of sensitive materials or even for helping to reduce the number of investigators ambushed and killed at crime scenes. By investigating these and other related directions, our aim is that opportunities could be found to contribute to a peaceful and happy society for all.
@@ Acknowledgements.
We gratefully acknowledge support from the Swedish Innovation Agency (Vinnova) for the project "AI-Powered Crime Scene Analysis" as well as advice from the Swedish Police Authority. The initial idea for this paper stemmed from a conversation with Mikael Lilja (Superintendent, Police Region Öst).References
@@
- Cooney et al. [2025] M. Cooney, S. Ponrajan, F. Alonso-Fernandez, Nano drone-based indoor crime scene analysis, arXiv preprint arXiv:2502.21019 (2025).
@@
- Esposito et al. [2023] M. Esposito, F. Sessa, G. Cocimano, P. Zuccarello, S. Roccuzzo, M. Salerno, Advances in technologies in crime scene investigation. diagnostics 2023, 13, 3169. h ps, doi. org/10.3390/diagnostics13203169 Academic Editor: Hiroshi Ikegaya Received 15 (2023).

### v2

@@ 3.1.2 Software
We also developed server and client programs to transmit LiDAR, RGB, and thermal data, as well as control signals between the gondola and our external laptop. TigerVNC was used to remotely develop the software on the Raspberry Pi.313131https://tigervnc.org/ As well, we built a user interface with color-coded controls in Python and OpenCV,323232https://opencv.org/ shown in Fig. 3. Various Arduino functions were implemented to allow the blimp to move, but to reduce risks if communication was lost, we ultimately chose to use commands that initiate short bursts rather than continuous motion.
A key challenge to our exploration was a lack of information on how to set up reasonably realistic mock-up crime scenes: e.g., what kind of evidence is typical and where it is usually found. We found no publicly available datasets of crime scene layouts—likely due to privacy, legal, or safety concerns, or regional variances on how to categorize crimes. Aside from two scans of mock-up crime scene shared by Galanakis et al. [32] and the Swedish Police from their National Forensics Center (NFC) training facility, and a few mock-up sketches on the internet, most search results seemed to point to story-writing tools (e.g., “Murder Scene Generator”).333333https://perchance.org/znevive9bn This was problematic, since, as the saying goes, "no two crime scenes are the same", blimps should be able to operate in different environments, and engineering evaluations typically require more than two or three trials (deep learning systems can even require thousands of training samples).
Thus, to support initial testing, we created two simplified simulation tools. The first lets users choose a crime type (e.g., homicide, assault, burglary, or arson) and floor plan (e.g., NFC Villa or our university’s 50 m2 "HINT" smart home), then randomly generates evidence types and locations, using customizable probabilities. The second tool places all evidence in a single heap, filling a grid by size—larger items first.
@@ 3.2 Exploring Possibilities for Indoor CSA Blimps
The developed platform was applied to check the questions identified in reviewing the literature: how much wind the blimp creates, how easy it is to control manually, how recordings of evidence can be further processed into 3D maps, and if thermal traces could be used. Looking ahead, we also aimed to explore the feasibility of automatic detection, classification of bloodstains, and autonomous path planning by the blimp for data acquisition. For evidence to use in testing, we gathered a photo of a firearm, two knives (one large and metallic, the other small and plastic), a pair of shoes, and three sheets with mock-up blood (red carmine dye—E120) to simulate passive, active, and transfer stains.
@@ 3.2.1 Disturbance
To test the degree to which sensitive evidence might be disturbed by a blimp, we placed a Mastech MS6252A digital anemometer on the ground beside the mock-up evidence. As a result, the wind sensor consistently read 0.0 m/s, even when the blimp passed close (20 cm) to the evidence. For comparison, flying a typical small drone—the DJI Ryze Tello—over the evidence at 1.2 m resulted in wind speeds of 0.6 to 0.8 m/s, which notably caused most of our evidence to scatter (viz., the sheets of paper containing mock-up blood and the photo of a firearm).
@@ 3.2.2 Sensing while Drifting
Next, we checked on the blimp’s ability to move and sense evidence. The blimp was manually controlled to obtain data ten times. In each trial, a random ordering of the evidence was calculated using our Python program, then the physical evidence was manually placed by a human. As a result, on average, the blimp captured 5.2 of 7 objects (74.3%) in its view. Some objects were occasionally missed because the blimp was susceptible to drafts and small imbalances that sometimes caused it to veer.
@@
For easy viewing, we explored converting video from the blimp to a 3D model. To test the waters, a video from the blimp was processed using the free version of 3DF Zephyr to create a 3D textured mesh.343434https://www.3dflow.net/3df-zephyr-photogrammetry-software/ The result is shown in Fig. 5(a). Some areas appeared rough—for example, around the mannequin "corpse" on the left side of the image—since the free software version only processes the first 50 frames extracted from the video. (Due to this limitation, we also sped up the video to ensure the entire scene was included.) However, the short time required to generate a 3D map (approximately 3 minutes) seemed promising.
@@ 3.2.4 Thermal Sensing
We also experimented with briefly touching a knife and the shoes before each trial controlling the blimp. As a result, some warm areas were visible in the thermal data, as exemplified in Fig. 5(b). However, in general, the heated objects—especially the smaller knife—seemed hard to distinguish due to the low resolution of our thermal sensor. As well, a checkerboard pattern appeared due to the blimp’s motion and the manner in which the MLX90640 sensor updates (in two passes, such that half of the pixels update slightly later). Thus, we believe that a higher-resolution sensor with a different update mechanism would be required to identify recently touched, non-metallic objects. That said, the MLX90640 could still be useful for other purposes, such as detecting investigators (to follow) or hidden persons.
@@ 3.2.5 Object Detection
Time and effort could be saved by automating key tasks such as detection of evidence. Detection results could also be highlighted in a 3D map in reports to investigators, or used to focus the blimp’s attention on areas requiring closer inspection. As an initial check, we tried passing an image taken by the blimp to a pretrained generic YOLOv8l detection model353535https://docs.ultralytics.com/ and ChatGPT.363636https://chatgpt.com As shown in Fig. 5(c), the YOLO model correctly identified one knife, but a bloody handprint was mistaken as a horse, and other objects were not detected. These errors were likely due to low image resolution or dim lighting conditions (approximately 300 lux in our lab when the image was taken). ChatGPT, by contrast, correctly detected all objects. This seemed promising and suggested the usefulness of applying a generic foundation model, if hardware permits and a specific model fine-tuned on a sufficient number of images of firearms and blood is unavailable.
@@ 3.2.6 Bloodstain Classification
To explore classifying the three main kinds of bloodstains, we tried processing a photo of our evidence "heap". First, we used color-picking and contours to identify clusters of red pixels. Then, we assessed the area and eccentricity of an ellipse fitted to each detected contour of reasonable size. We assumed that typical transfer stains (e.g., from a hand or foot) could be large compared to individual droplets, that passive drops should appear round (exhibit low eccentricity), and that active drops should appear elongated (have high eccentricity), as illustrated in Fig. 5(d). This resulted in a classification accuracy of 74.1%, with 20 of 27 detected blood contours correctly classified. Adding in false positives from color-picking (e.g., red tags on the shoes), the overall accuracy dropped to 69.0% (20 of 29), which still seemed promising for initial prototyping.
@@ 3.2.7 Coverage Path Planning for Crime Scenes
An automated blimp could operate without manual piloting. Given that a full solution to the complex problem of autonomous coverage path planning for a CSA blimp would not be possible within the scope of the current paper, we constrained our brainstorming here to identifying key variables, requirements and potential strategies. First, we propose that designs should consider the following factors:
Blimp’s capabilities
Movement precision: If the blimp’s movements are imprecise, it should fly at a safer, farther distance to avoid collisions and plan more overlap to ensure coverage.
Sensor accuracy: If sensors are less accurate, the blimp should fly closer to the evidence to capture sufficient detail.
Task
Task urgency: If time is constrained and the blimp is needed elsewhere, speed and efficiency can be prioritized. Else, in critical cases, more thorough scanning could be appropriate. For example, higher-detail documentation could require maneuvering beneath or around obstacles like tables and beds.
Task specification: If the blimp is expected to also gather evidence, it should move closer to objects of interest.
@@
Thus, in general, we believe that indoor crime scene mapping by a blimp should meet the following requirements, listed in arbitrary order:
Speed: Timely mapping is desired, prior to contamination or degradation of evidence, also given battery life constraints (e.g., mapping a room should take an hour or less).
Overlap: Some overlap (e.g., 25%) is required for 3D reconstruction, but excessive motion could disturb the scene.
Navigability: While some environments could be mostly empty and of simple rectangular shape, advanced capabilities could be required to navigate in complex domestic spaces with obstacles such as ceiling fans, lamps, chairs, or hallways, even if floor plans are known.
Efficiency: The blimp should minimize turning and vertical movement to conserve battery life. Flying at a greater height than the tallest standing object could be efficient, although indoor spaces constrain vertical movement more than outdoor ones.
Interaction: Humans, animals, or dynamic obstacles such as other robots could unintentionally or intentionally obstruct the blimp. Safety is paramount.
Adaptability: An advanced blimp could operate as part of a swarm, use an LLM to prioritize regions likely to contain evidence, and remain within a geofenced region to avoid disturbing people or objects.
Given these requirements, both simple and advanced planning approaches seemed possible:
Basic approaches: A snaking strategy combined with random walks, wall-following, or behavior-based robotics, at a collision-free height, could serve as a practical and intuitive baseline for early-stage testing.
Variable-height approaches: For maximum detail, the blimp could calculate a close path just above each surface. Conversely, initial high-altitude scans could be followed by targeted low-altitude re-visits to interesting areas (e.g., where potential evidence has been detected).
AI-driven approaches: Deep reinforcement learning (RL) could be used to mimic how investigators explore crime scenes. This could achieve high efficiency (minimizing flight time and avoiding redundant re-coverage), map quality (ensuring good feature overlap and loop closure, while avoiding poor viewing angles, low lighting, occlusions, shadows or reflective surfaces), room-agnosticism (functioning successfully in unknown or specially designed rooms or buildings, while avoiding hand-engineering edge cases), and robot-agnosticism (allowing different blimps to use the same process to learn how to move well). However, complex methods—such as deep RL—could also introduce challenges, including difficult implementation, large data requirements, extensive tuning requirements, and reduced explainability.
@@ 4 Discussion
In this paper, we explored a new design idea, of using a blimp to document sensitive evidence at an indoor crime scene.
Proof of Concept (Hardware and Software). To our knowledge, our new prototype is the first small indoor blimp to carry a thermal camera, three LiDARs, or a minicomputer (rather than a microcontroller). In addition to discussing various hardware challenges, some software was also developed to pilot the blimp and "generate" crime scenes.
Manual operation. As a result of our exploration, we found some further support that drones can disturb sensitive evidence and confirmed our premise that blimps can keep wind to a minimum. As well, manually piloting was possible but imperfect, with an average of 74.3% of objects captured in view in a single pass. Furthermore, a 3D map was generated within a few minutes using free software, and some recommendations for thermal sensing were given.
Towards Automation. All objects in the blimp’s camera view were detected correctly by ChatGPT. As well, the three main kinds of bloodstain were detected and classified with an accuracy of 74.1% via color-picking and consideration of contour area and eccentricity. Finally, we discussed path-planning considerations and strategies.
As well, a video and code have been made available.373737https://youtu.be/5ep9UeChn68,383838https://github.com/martincooney/Crime_Blimp
@@ 4.1 Limitations and Future Work
The current study is limited by its exploratory design approach, working with mock-up evidence in a lab:
Unoptimized components and guesswork. To reduce size and weight, surface mount or low power components (e.g., Raspberry Pi Pico instead of Zero 2 W), Dist-YOLO in place of LiDARs [33], or smaller batteries could be evaluated. Hybrid setups could also allow greater speed and payloads. As well, our current reliance on a remote laptop—though common in related work—would pose security risks at a real crime scene, given that criminals could find a way to hack (intercept or jam) communications. Thus, dedicated onboard hardware acceleration for computer vision is desired. For basic tasks, an NPU (neural processing unit) could suffice, such as the 1 GHz unit on OpenMV N6 that can run YOLO at 30 FPS.393939https://openmv.io/collections/all-products/products/openmv-n6 Heavier Foundation Models or LLMs requiring GPUs (Graphic Processor Units) with sufficient VRAM (Video Random Access Memory) could also become feasible on future blimp platforms; e.g., 16 GB for a 30B LLM with 4-16 bit quantization.404040https://pyvideotrans.com/en/_posts/deployllm Our crime scene generation software is also heuristic-based and would benefit from training with real data to enable more realistic outputs.
Simplified approaches. Control algorithms should be developed to help blimp operators. Given the low thrust of blimp propellers, we should clarify if drifting might not even be required in some contexts. Additional use cases for thermal sensing and alternative miniature thermal cameras should be explored, as well as other free 3D Mapping tools (that allow more than fifty photos). Operation in challenging conditions—such as darkness or smoke—should also be explored, for example using auxiliary lighting, low-light cameras, or near-infrared imaging. Ultimately, field evaluations with real crime scenes and data will be essential to develop real products and services.
Limited data and testing. A more comprehensive study is required to identify the various kinds of evidence that could be important and how they can be detected. Also, for classifying bloodstains, only a single image was used; a reasonably-sized dataset—ideally real—should be used to train and evaluate a more advanced machine learning model, to achieve better results. Furthermore, in-depth forensic interpretation of bloodstain composites could reveal valuable information such as weapon type, handedness, number of blows, sequence of injuries, sources of transfer stains (e.g., a hand or a shoe), or immediacy of death. Path planning should also be implemented and tested; for this, an omnidirectional camera could also improve scene coverage while minimizing unnecessary flight, reducing the risk of disturbing sensitive evidence.
Beyond documentation of evidence at indoor crime scenes, remote-controlled blimps could even aid in hostage or standoff situations by offering silent surveillance to inform response planning, and reduce investigator exposure to ambushes or danger. Blimps could also benefit other sensitive applications where airflow or noise from conventional drones could be detrimental—for instance, munitions or chemical facilities with ignition risks, nuclear disaster sites where radioactive dust should not be stirred, avalanche or building sites in danger of collapsing, habitats of endangered animals sensitive to noise, or pharmaceutical cleanrooms or semiconductor plants where slight air movements could introduce flaws. By pursuing such directions, our aim is to identify opportunities for technology to contribute to a more peaceful and happy society for all.
@@ Acknowledgements.
We gratefully acknowledge support from the Swedish Innovation Agency (Vinnova) for the project "AI-Powered Crime Scene Analysis" as well as various advice from the Swedish Police Authority (the initial idea for this paper stemmed from a conversation with Mikael Lilja).References
- un2 [2023] The Sustainable Development Goals Report 2023, 2023. URL: https://unstats.un.org/sdgs/report/2023/The-Sustainable-Development-Goals-Report-2023.pdf, accessed: 2025-05-26.
@@
- Cooney et al. [2025] M. Cooney, S. Ponrajan, F. Alonso-Fernandez, Nano Drone-based Indoor Crime Scene Analysis, 2025 IEEE International Conference on Advanced Robotics and its Social Impacts (ARSO) (accepted April 29, 2025; arXiv preprint arXiv:2502.21019) (2025).
@@
- Hong and Tanaka [2025] D. Hong, Y. Tanaka, Buoyant choreographies: Harmonies of light, sound, and human connection, in: IEEE International Conference on Robotics and Automation (ICRA) 2025, IEEE, 2025.
@@
- Esposito et al. [2023] M. Esposito, F. Sessa, G. Cocimano, P. Zuccarello, S. Roccuzzo, M. Salerno, Advances in Technologies in Crime Scene Investigation, Diagnostics 2023, 13(20), 3169. DOI org/10.3390/diagnostics13203169 (2023).